Derek will deposit $3,227.00 per year for 14.00 years into an account that earns 9.00%. Assuming the first deposit is made 7.00 years from today, how much will be in the account 36.00 years from today?

What is the value today of receiving $2,664.00 per year forever? Assume the first payment is made next year and the discount rate is 13.00%.
Answer #1

1.Present value=Cash flows*Present value of discounting factor(rate%,time period)

=3227/1.09^7+3227/1.09^8+...............+3227/1.09^20

=3227[1/1.09^7+1/1.09^8+...............+1/1.09^20 ]

=3227*4.642627079

=$14981.76(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A =$14981.76*(1.09)^36

=$333,362.46(Approx).

2.Present value of perpetuity=Annual cash flows/discount rate

=2664/0.13

=$20492.31(Approx).